Asilekpe Company Limited is considering constructing a borehole in the village of Adukofe. The 

construction would cost Gh₵2 million and there will be no cost of maintenance. The following

table shows the company’s anticipated demand over the lifetime of the bridge:

Price per gallon (Gh₵) Number of Gallons (‘000)

8 0

7 100

6 200

5 300

4 400

3 500

2 600

1 700

0 800

a. If the company were to construct the borehole what would be its profit-maximizing 

price? Would that be the efficient level of output? Explain your answer


1
Expert's answer
2021-03-16T08:52:31-0400

(a) "4*400000=1600000"

"Gh\u20b5 1,600,000" million

(b)No. The company is making a loss of four hundred thousand

"=2000000-1600000=400000"